OpenStudy (anonymous):

Rationalize the denominator: \[\frac{1}{((x)^(1/3))-((y)^(1/3))}\]

OpenStudy (anonymous):

\[\frac{1}{\sqrt[3]{x}-\sqrt[3]{y}}\]

OpenStudy (anonymous):

Change the sign first for bottom one than multiply & divide by the same to current..than solve it.

OpenStudy (anonymous):

I did that and I got it wrong...

OpenStudy (anonymous):

okey let me try..